# BLACKOUT: Data-Oblivious Computation with Blinded Capabilities

### v3

@@ 5.3. Motivating Examples
The compile-time analysis is dependent on the developer providing (initial) information about which variables are expected to contain sensitive information through blinded attributes (__blinded, ❷ in LABEL:lst:oblivious_select). Optionally, the developer can also declare functions blinded (__attribute__((blinded)), ❶ in LABEL:lst:oblivious_select). This clearly identifies functions expected to return blinded data, improving developer ergonomics. The compile-time analysis is not limited to the information provided by the developer. For example, in LABEL:lst:oblivious_select ❸, the compiler can infer the variable res must be blinded due to the assignment from c tainting it at ❺. However, annotating functions as blinded improves the readability of data-oblivious code.
@@ 6.1. BLACKOUT CHERI-RISC-V CPU
We implement blinded capabilities in register-transfer level (RTL) on CHERI-Toooba (Fuchs et al., 2024), a CHERI-enabled RISC-V processor based on Bluespec’s speculative out-of-order Toooba IP (Nikhil, 2025). Unlike previous data-oblivious ISAs (Yu et al., 2019; ElAtali et al., 2024b), BLACKOUT does not add any additional instructions to the underlying ISA; all necessary architectural changes are achieved by adjusting the CHERI permission model.
@@
We prohibit valid capabilities from being blinded (I2, § 5). The execution of any capability-modifying instruction with operands that would result in a blinded capability causes a fault. Enforcing this invariant allows us to avoid unnecessarily making capability-modifying instructions data-oblivious. Note that this enforcement does not break the CHERI programming model; “blinded blinded capabilities” are themselves redundant as they can never be dereferenced to access memory in order to maintain confidentiality.
@@
However, another challenge arises: the non-blinded csp generates non-blinded data when storing spilled values. This causes ambiguity when restoring spilled registers, as the CPU cannot distinguish between blinded and non-blinded register spills. To address this, BLACKOUT’s csc emits special blinded register records (BRRs) as result spilling blinded registers. A BRR is a 128-bit structure containing the spilled value and 64-bit marker. This marker differentiates BRRs from conventional CHERI capabilities and maintains their validity tags. When spilled capability register is restored, the validity tag signals that the value restored from memory is either a capability or a BRR. The marker indicates whether the value is a BRR, ensuring the destination register is correctly blinded.
@@
As mentioned in § 2.2, hardware optimizations, such as speculation, can inadvertently introduce side channels when operating on secrets. To protect blinded data against such leakage, we ensure that any blinded data flowing to decision-making parts of the hardware (such as branch predictors) are zeroed out. This guarantees that execution is truly oblivious to blinded data, both architecturally and transiently. This is similar to approaches used in prior work (Yu et al., 2019; ElAtali et al., 2024b).
@@ 6.2. BLACKOUT Software Stack
We extended the Cornucopia revocation mechanism to provide a blinded_malloc API that returns blinded capabilities to blinded heap allocations. To meet I3 (§ 5), such blinded heap allocations must not overlap with other, non-blinded allocation. This ensures the blinded capability returned blinded_malloc has exclusive access to the newly created allocation. The blinded_malloc API relies on Cornucopia to ensure the blinded capability does not overlap with any concurrently existing capabilities. BLACKOUT’s blinded_malloc is integrated into CheriBSD via the MRS (similar to the integration of Cornucopia, discussed in § 2.1). This makes blinded capabilities allocator-agnostic, allowing userspace processes to employ different underlying allocators as long as allocations occur via the shim.
@@
Recall from § 4.2 that blinded memory must be securely reclaimed to prevent information leaking from previously blinded memory regions. BLACKOUT implements explicit zeroing policies for blinded heap and stack allocations. For blinded heap allocation, the free API, implemented in the shim, inspects whether memory to be deallocated is blinded and erases the contents of blinded allocations before freeing it. For stack-allocated blinded variables, the compiler automatically inserts a memset IR instruction to zero out their memory immediately after the variable’s lifetime ends. This ensures sensitive data from blinded variables is securely erased, preventing unintended reuse or leakage when stack frames are reallocated.
@@
As discussed in § 6.1, any newly created capability must initially be configured with its “non-blinded” permission bit set. To support blinded capabilities in CheriBSD, we perform a thorough scan of all CheriBSD code to ensure that any derived permissions (e.g., for drivers or userspace) also unset the “non-blinded” permissions. We also extend the CheriBSD CPU exception handler to handle Blinded capability exceptions, and adapted the signal code with the newly introduced from the Toooba Core. We integrated blinded capabilities into CheriBSD 24.05, resulting in a total of 100 lines of code changes over 14 distinct files.
@@ 7.1. Power & resource usage
Blinded Cheri Toooba
@@ 7.2. Performance
Second, for blinded workloads, we evaluate the performance impact of BLACKOUT using five OISA benchmarks adapted from Yu et al. (Yu et al., 2019). To adapt them to BLACKOUT, we simply blind secret inputs with the __blinded attribute, change dynamic allocations to use our blinded allocator API, and compile the benchmarks using our blinded capability-enhanced compiler. We measured the performance on CheriBSD with BLACKOUT support in pure-capability mode with the Cornucopia revocation mechanism enabled. Our setup follows the CheriBSD benchmark guide (Digital Security by Design, 2024), with the exception of running the BLACKOUT benchmarks in pure-capability mode (§ 2.1) with blinded capabilities, and enabling the experimental Cornucopia revocation mechanism, which is essential to guarantee exclusive access for blinded capabilities. We compiled all benchmarks with -O3 and measured the combined user and system time using the time command. This measures the entire lifetime of blinded capabilities—including capability creation, algorithm execution, and memory reclamation (i.e., zeroing out blinded memory). For the findmax, binary_search, and integer_sort benchmarks, we used blinded input arrays containing $2^{N}$ integers. For the matrix multiplication benchmark, we multiplied two square matrices of size $2^{\frac{N}{2}}\times 2^{\frac{N}{2}}$. In the DNN example, the blinded input size was $2^{\frac{N}{2}}$, and the network consisted of two layers, each of size $2^{\frac{N}{2}}\times 2^{\frac{N}{2}}$, with a fixed output size of $2^{6}$. We evaluated all benchmarks with $N\in\{12,14,16,18,20\}.$
@@
The geometric mean overhead per benchmark is shown in Figure 4(f). The overall result shows a minimal geometric mean overhead of $1.5\%$ for BLACKOUT compared to the purecap configuration, which is several times lower than overheads for prior work enforcing data-oblivious computation (Yu et al., 2019; ElAtali et al., 2024b). The overhead in blinded workloads is caused by several factors:
@@
    Clearing blinded memory at the end of a blinded capability’s lifecycle. This is done by the compiler for blinded memory on the stack on function returns, and by the blinded_malloc for blinded memory on the heap on freeing. For instance, when N = 20, binary search operates on a 4MiB array. In this setting, memory reclamation accounts (zeroing out) is attributed to the most of the overhead.
@@
[…]
@@ 7.3. Security
    Standard CHERI-provided memory safety. This includes attacks based on memory safety violations, such as out-of-bounds access and use-after-free. We do not loosen any of the restrictions imposed by standard CHERI (such as proper capability bounds and lifetime). We are therefore compliant with the standard CHERI model, and as such inherit its safety guarantees against spatial and temporal memory violations. This extends to the guarantees provided by formal models (Gao and Melham, 2021; Antón et al., 2024; Ploix et al., 2025).
@@
    Side-channel protection for blinded data. This is provided by the enforcement of data-oblivious computation on blinded data (I5 and I4). Any violation of this results in a fault, as explained in § 5. We inherit formal guarantees for this from the OISA and BliMe models (Yu et al., 2019; ElAtali et al., 2024b). We further show its efficacy empirically by running the OISA benchmarks (Yu et al., 2019) using BLACKOUT and verify that introducing non-data-oblivious changes to them is either detected by our compiler or causes a run-time fault.
@@
    Confidentiality of blinded data with respect to direct access. While BliMe guarantees the confidentiality of blinded data in memory using in-memory tags, we achieve the same goal by ensuring that 1) blinded capabilities have exclusive access to blinded data throughout their lifetime (I3), 2) memory containing blinded data is cleared at the end of the corresponding blinded capability’s lifetime (also I3), and 3) blinded data cannot be stored to memory using unblinded capabilities (I1). The resulting combination ensures that, as in OISA and BliMe, any data stored as blinded into memory, is loaded as blinded into registers. This holds until the data is explicitly declassified as discussed in § 5.
@@ 8. Discussion & Limitations
As discussed in § 5, BLACKOUT ensures memory designated as blinded (either annotated by the developer, or inferred as such by the compiler) is made blinded at allocation time. This ensures exclusive access through corresponding blinded capabilities, satisfying invariants I1, I2 and I3 (§ 5). However, some correct data-oblivious programs may initially allocate memory as non-blinded, becoming blinded only after interactions with other blinded data. While BLACKOUT restricts this pattern, a less-restrictive variation could allow dynamically blinding previously allocated memory. This would involve scanning program memory to revoke existing, non-blinded capabilities referencing soon-to-be-blinded memory, similar to revocation methods proposed by Cornucopia (Wesley Filardo et al., 2020) and Cornucopia Reloaded (Filardo et al., 2024)). The drawback of this approach is the performance overhead due to memory scanning. Future work might explore dynamic revocation methods and performance optimization techniques, such as delayed scanning, to minimize immediate overhead while ensuring eventual exclusive access.
Similarly, data-oblivious software may eventually need to unblind memory once data-oblivious operations on the data conclude and intermediate secret values have been cleared. Concurrent data-oblivious operations on shared blinded memory could necessitate revoking overlapping blinded capabilities to maintain temporal memory safety, preventing unauthorized reuse after unblinding. Currently, BLACKOUT does not address this scenario of oblivious-data-race-safety, leaving it for future exploration.
@@
Another solution for unblinding blinded memory containing non-sensitive results is leveraging CHERI’s facilities for capability sealing (§ 2.1). Sealed capabilities help to protect against tampering or unauthorized modification of capabilities. Once a capability is sealed, its properties, such as permissions and bounds, are fixed, and any attempt to tamper with the capability will result in invalidating it. In the context of blinded memory, sealed capabilities can be used to create a “capability escrow” which holds a non-blinded version of a blinded capability for safe-keeping until the data-oblivious operations have concluded and the results needs to be read out of blinded memory. To achieve this, software can use a blinded capability to seal its non-blinded counterpart. In the sealed state, the non-blinded capability cannot be dereferenced and thus poses no threat to the confidentiality of blinded memory. Once the computation has completed, any remaining secret intermediate values should be cleared, and the blinded capability used to unseal its counterpart, allowing access to the result. We leave the exploration of software APIs for capability escrow for future work.
@@
In BLACKOUT, the developer uses __blinded annotations to inform the compiler that a variable is blinded (§ 6.2). Compiler attributes are a common way for compiler developers to allow additional information about program semantics to be communicated to the compiler without changing the underlying language. This approach also avoids the need for intrusive changes to the compiler pipeline as it is implemented through Clang’s existing annotate_type attribute. A limitation of this approach is that introducing annotations for e.g., positional parameters in functions is cumbersome within the confines of existing types of annotations. Another approach would be to introduce a new “blinded C” dialect that integrates blindedness into the type system (as discussed for Rust in § 1). Blinded C could incorporate a blinded attribute keyword similar to the const specifier, which would turn a variable into a blinded counterpart of its basic type. While this can be more ergonomic for the developer, allowing function parameters to be annotated easily to enhance the compiler analysis, it requires intrusive changes to the compiler and language and comes at the cost of maintainability and extensibility, which could be burdensome for architectures and software stacks that are still evolving, such as CHERI.
@@
As we note in § 6.1, BLACKOUT, unlike previous data-oblivious ISAs (Yu et al., 2019; ElAtali et al., 2024b), does not add any additional instructions to the underlying ISA but leverages existing CHERI functionality with only small adjustments to its permission model to facilitate necessary architectural changes to support blinded capabilities. A line of future work is to extend the ISA by adding dedicated instructions to support blinded memory. A potential extension is an instruction to directly set the blindedness bit in a register without a load via a blinded capability. This would allow the blinded capability-enhanced compiler to optimize code where a certain variable can be kept completely in a register throughout its lifetime. Currently, BLACKOUT requires such variables to be allocated on the stack in order to load them via the corresponding blinded capability.
@@
BLACKOUT demonstrates a practical method for enhancing CHERI’s security model without invasive architectural changes by integrating data-oblivious computation capabilities into the existing CHERI protection model. This integration provides insights into how capability-based architectures can evolve to support advanced security properties, such as data-oblivious computation, while maintaining compatibility and performance. Consequently, BLACKOUT’s principles and mechanisms can inform future discussions and proposals in ongoing CHERI standardization efforts (CheriAlliance, 2025), potentially guiding the evolution of CHERI-enabled architectures towards broader security guarantees. Above, we discussed adding custom instructions. But maintaining the CHERI ISA as we currently do in BLACKOUT provides backward-compatibility for non-BLACKOUT hardware and can ease integration into the standard.

### v4

@@ 5.3. Motivating Examples
The compile-time analysis is dependent on the developer providing (initial) information about which variables are expected to contain sensitive information through blinded attributes (__blinded, ❷ in LABEL:lst:oblivious_select). Optionally, the developer can also declare functions blinded (__attribute__((blinded)), ❶ in LABEL:lst:oblivious_select). This clearly identifies functions expected to return blinded data, improving developer ergonomics by enhancing the readability of data-oblivious code. The compile-time analysis is not limited to the information provided by the developer. For example, in LABEL:lst:oblivious_select ❸, the compiler can infer the variable res must be blinded due to the assignment from c tainting it at ❺.
@@ 6.1. BLACKOUT CHERI-RISC-V CPU
We implement blinded capabilities in register-transfer level (RTL) on CHERI-Toooba (Fuchs et al., 2024), a CHERI-enabled RISC-V processor based on Bluespec’s speculative out-of-order Toooba IP (Nikhil, 2025). Unlike previous data-oblivious ISAs (Yu et al., 2019a; ElAtali et al., 2024b), BLACKOUT does not add any additional instructions to the underlying ISA; all architectural changes are achieved by adjusting the CHERI permission model.
@@
We prohibit valid capabilities from being blinded (I2, § 5). The execution of any capability-modifying instruction with operands that would result in a blinded capability causes a fault. Enforcing this invariant allows us to avoid unnecessarily making capability-modifying instructions data-oblivious. Note that this enforcement does not break the CHERI programming model; “blinded blinded capabilities” are themselves redundant as they can never be dereferenced to access memory in order to maintain confidentiality (I5, § 5).
@@
However, another challenge arises: the non-blinded csp generates non-blinded data when storing spilled values. This causes ambiguity when restoring spilled registers, as the CPU cannot distinguish between blinded and non-blinded register spills. To address this, BLACKOUT’s csc emits special blinded register records (BRRs) as a result of spilling blinded registers. A BRR is a 128-bit structure containing the spilled value and 64-bit marker. This marker differentiates BRRs from conventional CHERI capabilities and maintains their validity tags. When a spilled capability register is restored, the validity tag signals that the value restored from memory is either a capability or a BRR. The marker indicates whether the value is a BRR, ensuring the destination register is correctly blinded.
@@
As mentioned in § 2.2, hardware optimizations, such as speculation, can inadvertently introduce side channels when operating on secrets. To protect blinded data against such leakage, we ensure that any blinded data flowing to decision-making parts of the hardware (such as branch predictors) are zeroed out. This guarantees that execution is truly oblivious to blinded data, both architecturally and transiently. This is similar to approaches used in prior work (Yu et al., 2019a; ElAtali et al., 2024b).
@@ 6.2. BLACKOUT Software Stack
We extended the Cornucopia revocation mechanism to provide a blinded_malloc API that returns blinded capabilities to blinded heap allocations. To meet I3 (§ 5), such blinded heap allocations must not overlap with other, non-blinded allocations. This ensures the blinded capability returned blinded_malloc has exclusive access to the newly created allocation. The blinded_malloc API relies on Cornucopia to ensure the blinded capability does not overlap with any concurrently existing capabilities. BLACKOUT’s blinded_malloc is integrated into CheriBSD via the MRS (similar to the integration of Cornucopia, discussed in § 2.1). This makes blinded capabilities allocator-agnostic, allowing userspace processes to employ different underlying allocators as long as allocations occur via the shim.
@@
Recall from § 4.2 that blinded memory must be securely reclaimed to prevent information leaking from previously blinded memory regions. BLACKOUT implements explicit zeroing policies for blinded heap and stack allocations. For blinded heap allocation, the free API, implemented in the shim, inspects whether memory to be deallocated is blinded and erases the contents of blinded allocations before freeing it. For stack-allocated blinded variables, the compiler automatically inserts a memset IR instruction to zero out their memory immediately after the variable’s lifetime ends. This ensures sensitive data from blinded variables is securely erased, preventing unintended reuse or leakage when stack frames are reallocated. Blinded global data persists until the process terminates and consequently does not need to be reclaimed. The contents of (physical) pages are zeroed out by the OS before they are recycled for other processes.
@@
As discussed in § 6.1, any newly created capability must initially be configured with its “non-blinded” permission bit set. To support blinded capabilities in CheriBSD, we perform a thorough scan of all CheriBSD code to ensure that any derived permissions (e.g., for drivers or userspace) also unset the “non-blinded” permissions. We also extend the CheriBSD CPU exception and signal handlers to handle blinded capability exceptions and the new signal introduced to the BLACKOUT CHERI-Toooba Core. We integrated blinded capabilities into CheriBSD 24.05, resulting in a total of 100 lines of code changes over 14 distinct files.
@@ 7.1. Power & resource usage
Blinded CHERI-Toooba
@@ 7.2. Performance
OISA benchmarks.
Second, for blinded workloads, we evaluate the performance impact of BLACKOUT using five OISA benchmarks adapted from Yu et al. (Yu et al., 2019a). To adapt them to BLACKOUT, we simply blind secret inputs with the __blinded attribute, change dynamic allocations to use our blinded allocator API, and compile the benchmarks using our blinded capability-enhanced compiler. Porting the OISA benchmarks to BLACKOUT took 1-5 LoC changes (¡1%) per benchmark. We measured the performance on CheriBSD with BLACKOUT support in pure-capability mode with the Cornucopia revocation mechanism enabled. Our setup follows the CheriBSD benchmark guide (Digital Security by Design, 2024), with the exception of running the BLACKOUT benchmarks in pure-capability mode (§ 2.1) with blinded capabilities, and enabling the experimental Cornucopia revocation mechanism, which is essential to guarantee exclusive access for blinded capabilities. We compiled all benchmarks with -O3 and measured the combined user and system time using the time command. This measures the entire lifetime of blinded capabilities—including capability creation, algorithm execution, and memory reclamation (i.e., zeroing out blinded memory). For the findmax, binary_search, and integer_sort benchmarks, we used blinded input arrays containing $2^{N}$ integers. For the matrix multiplication benchmark, we multiplied two square matrices of size $2^{\frac{N}{2}}\times 2^{\frac{N}{2}}$. In the DNN example, the blinded input size was $2^{\frac{N}{2}}$, and the network consisted of two layers, each of size $2^{\frac{N}{2}}\times 2^{\frac{N}{2}}$, with a fixed output size of $2^{6}$. We evaluated all benchmarks with $N\in\{12,14,16,18,20\}.$
@@
The geometric mean overhead per benchmark is shown in Figure 4(f). The overall result shows a minimal geometric mean overhead of $1.5\%$ for BLACKOUT compared to the purecap configuration, which is several times lower than overheads for prior work enforcing data-oblivious computation (Yu et al., 2019a; ElAtali et al., 2024b). The overhead in blinded workloads is caused by several factors:
@@
    Clearing blinded memory at the end of a blinded capability’s lifecycle. This is done by the compiler for blinded memory on the stack on function returns, and by the blinded_malloc for blinded memory on the heap on freeing. For instance, when N = 20, binary search operates on a 4MiB array. In this setting, memory reclamation (zeroing) accounts for most of the overhead.
@@
We also evaluate the combined performance impact of CHERI’s memory-safety and BLACKOUT’s data-oblivious enforcement. We measure a moderate geometric mean overhead of $23.5\%$ compared to the unprotected baseline configuration with capability-enforcement completely disabled. We observe that the overhead for both CHERI’s purecap and BLACKOUT’s purecap + blinded modes relative to the unprotected baseline is significantly larger in the binary_search benchmark, particularly for smaller input sizes, compared to the other OISA benchmarks. We investigate further the composition of the overhead in that benchmark. In consultation with the University of Cambridge Computer Laboratory Security Group we identified three possible reasons for the poor performance:
@@
SpectreGuard benchmarks.
Third, we run the SpectreGuard (Fustos et al., 2019) synthetic benchmark using the version adapted by Daniel et al. (Daniel et al., 2023) to evaluate ProSpeCT (see § 9). This version uses data-oblivious implementations of the chacha20, sha2, and curve25519 cryptographic algorithms from HACL*(Zinzindohoué et al., 2017), a formally verified cryptographic library written in F* (Swamy et al., 2016). The benchmarks represent a workload consisting of public-data computations, which gain substantial performance from speculative execution, alongside encryption routines, which are less impacted by speculation. Each benchmark varies the proportion of speculative versus cryptographic work (S/C), as shown in Table 4. The ProSpect benchmark version already annotates all secret values in those test cases. We adapt those annotations to blinded attributes and run all the test cases on BLACKOUT CHERI-Toooba in both purecap and purecap + blinded versions. Table 4 shows the results (average of 20 runs) for the chacha20 and sha2 test cases. Unlike for the OISA benchmark results, where BLACKOUT’s overhead could be attributed to zeroing out stack and heap variables, the ProSpeCT version of the SpectreGuard benchmark places all secret variables in static variables allocated from the program’s data section, similar to global variables. Consequently, BLACKOUT does not introduce any discernible overhead for encryption time or workload time. The slight speed improvement of the purecap + blinded version compared to the purecap baseline in chacha20 25S/7C and sha2 90S/10C falls within standard deviation (maximum $\sigma=0.08\%$). For completeness, we also include results for nocap, which performs worse than purecap in most cases. We attribute this difference to improvements in the RISC-V backend for CHERI-RISC-V targets compared to plain RISC-V targets in CHERI-LLVM, resulting in more compact and efficient assembly. We confirmed this through manual investigation.
For the Curve25519 test case, we discovered a constant-time violation caused by a secret-dependent branch instruction in the compiled binary. Although this violation is not present in the formally verified source code, it is introduced by the compiler, which in our case is CHERI’s fork of Clang. This violation is not reported in ProSpeCT’s evaluation because it uses the GCC compiler, which likely correctly avoided generating such a secret-dependent branch. This difference in generated assembly code highlights the usefulness of BLACKOUT: different compilers and/or configurations can silently introduce constant-time violations which BLACKOUT can detect and stop even if the source does not contain such a violation.
Table 4. SpectreGuard benchmark performance average measured over 20 runs of each test case (maximum $\sigma=0.08\%$).
Blinded CHERI-Toooba 25S/75C 50S/50C 75S/25C 90S/10C
Total Ticks $\Delta$ (to row above) Total Ticks $\Delta$ (to row above) Total Ticks $\Delta$ (to row above) Total Ticks $\Delta$ (to row above)
chacha20
nocap 25860874 23132682 25417083 20661078
purecap 22546871 -3314003 ($12.81\%$) 20378637 -2754045 ($11.91\%$) 25430015 12932 ($0.05\%$) 20256607 -404472 ($1.96\%$)
purecap + blinded 22541024 -5847 ($0.03\%$) 20378935 298 (0.00%) 25434929 4914 ($0.02\%$) 20256061 -546 ($0.00\%$)
sha2
nocap 24819790 21595015 25417599 22163550
purecap 24771185 -48604 ($0.20\%$) 21591164 -3851 ($0.02\%$) 26342237 924638 ($3.64\%$) 22378912 215362 ($0.97\%$)
purecap + blinded 24771730 544 ($0.00\%$) 21593499 2335 ($0.01\%$) 26342910 672 ($0.00\%$) 22371464 -7448 ($0.03\%$)
@@ 7.3. Security
Empirical Spectre mitigation evaluation.
CHERI-Toooba is vulnerable to Spectre branch target buffer (BTB) (Kocher et al., 2019), return stack buffer (RSB) (Maisuradze and Rossow, 2018; Koruyeh et al., 2018) and store to load (STL) (Horn, 2018) (cf. test suite in (Fuchs et al., 2021)). We blinded the secret value in all test cases by inserting a single candperm instruction. As seen in Table 5, BLACKOUT prevents all Spectre attacks from (Fuchs et al., 2021) because BLACKOUT catches side-channel violations even during speculation, but suppresses faults until speculation is confirmed (and ignores them otherwise). Additionally, we have replicated ProSpeCT’s (Daniel et al., 2023) Spectre tests. However, as they involve dereferencing a secret data value as a pointer, they are inherently prevented by CHERI’s architectural security properties, which preclude dereferencing non-capability data. Note that none of the Spectre tests in (Fuchs et al., 2021; Daniel et al., 2023) use transient capability forgery, and therefore do not require CSC to prevent.
Table 5. Transient-execution attacks successfully prevented on CHERI-Toooba and Blinded CHERI-Toooba. ✓ indicates the attack is successfully defended against, while ✗ indicates the attack succeeds.
Spectre variant
PHT BTB RSB STL
CHERI-Toooba
Blinded CHERI-Toooba
Empirical non-interference evaluation.
Data-oblivious software inherently provides the non-interference property (Winderix et al., 2024a). We verify this empirically using the methodology from (Winderix et al., 2024a, b; Bognar et al., 2023): relevant signal traces are extracted from the value change dump (VCD) waveforms across different program runs to ensure that runs with different secret data values produce identical signal traces. As in prior work (Winderix et al., 2024a), we extract traces for signals that could leak secret data through cache-timing (addresses of cache accesses), speculative execution (branch predictor states), and port contention and instruction latencies (reorder buffer scheduling). As BLACKOUT enforces data-oblivious processing of blinded data, we use the data-oblivious binary_search program from (Yu et al., 2019a) as a case study. We vary the blinded data values between two runs of the program, and verify that the extracted traces for the relevant signals are identical.
Theoretical security evaluation.
@@
    Standard CHERI-provided memory safety. This includes attacks based on memory safety violations, such as out-of-bounds access and use-after-free. We do not loosen any of the restrictions imposed by standard CHERI (such as proper capability bounds and lifetime). We are therefore compliant with the standard CHERI model, and as such inherit its safety guarantees against spatial and temporal memory violations. This extends to the guarantees provided by formal models (Gao and Melham, 2021; Duque Antón et al., 2025; Ploix et al., 2025).
@@
    Side-channel protection for blinded data. This is provided by the enforcement of data-oblivious computation on blinded data (I5 and I4). Any violation of this results in a fault, as explained in § 5. We inherit formal guarantees for this from the OISA and BliMe models (Yu et al., 2019a; ElAtali et al., 2024b). We further show its efficacy empirically by running the OISA benchmarks (Yu et al., 2019a) using BLACKOUT and verify that introducing non-data-oblivious changes to them is either detected by our compiler or causes a run-time fault.
@@
    Confidentiality of blinded data with respect to direct access. While BliMe guarantees the confidentiality of blinded data in memory using in-memory tags, we achieve the same goal by ensuring that 1) blinded capabilities have exclusive access to blinded data throughout their lifetime (I3), 2) memory containing blinded data is cleared at the end of the corresponding blinded capability’s lifetime (also I3), and 3) blinded data cannot be stored to memory using unblinded capabilities (I1). The resulting combination ensures that, as in OISA and BliMe, any data stored as blinded into memory, is loaded as blinded into registers. This holds until the data is explicitly declassified (§ 5).
@@ 8. Discussion & Limitations
As discussed in § 5, BLACKOUT ensures that memory designated as blinded—either annotated by the developer, or inferred as such by the compiler—is blinded at allocation time. This ensures exclusive access through corresponding blinded capabilities, satisfying invariants I1, I2 and I3 (§ 5). However, some correct data-oblivious programs may initially allocate memory as non-blinded, only requiring it to be blinded after interacting with other blinded data. While BLACKOUT restricts this pattern, a less-restrictive variation could allow dynamically blinding previously allocated memory. This would require scanning program memory to revoke existing, non-blinded capabilities referencing to-be-blinded memory, similar to revocation strategies used by Cornucopia (Wesley Filardo et al., 2020) and Cornucopia Reloaded (Filardo et al., 2024). The trade-off is increased performance overhead. Future work might explore dynamic or delayed revocation strategies to reduce this cost while ensuring eventual exclusive access.
Eventually, data-oblivious software must unblind memory after data-oblivious computation concludes and secret intermediates are cleared (§ 5). In concurrent scenarios, shared blinded memory may require revoking overlapping blinded capabilities to maintain temporal memory safety and prevent unauthorized reuse after unblinding. Currently, BLACKOUT does not address this oblivious-data-race-safety, leaving it as an open problem for future work.
@@
A solution for further securing the unblinding of blinded memory containing non-sensitive results is to leverage CHERI’s facilities for capability sealing (§ 2.1). Sealed capabilities cannot be de-referenced and protect against tampering by fixing properties like permissions and bounds; any attempt to tamper with the capability will result in invalidating it. In this context, sealed capabilities enable a form of “capability escrow”, where a non-blinded version of a blinded capability stored for safe-keeping until the data-oblivious operations complete. Software can use a blinded capability to seal its non-blinded counterpart, which cannot be de-referenced while sealed and thus poses no threat to confidentiality. After data-oblivious computation has completed, secret intermediate values are cleared, and the blinded capability is used to unseal the non-blinded counterpart, allowing access to the result. Designing software APIs for capability escrow is left for future work.
@@
In BLACKOUT, developers use __blinded annotations to indicate blinded variables (§ 6.2). Compiler attributes are a common mechanism for conveying semantic information without changing the underlying language. BLACKOUT leverages Clang’s existing annotate_type mechanism, available since LLVM 14, avoiding intrusive compiler changes.
However, this approach has limitations—annotating positional parameters in functions can be cumbersome within the confines of existing types of annotations. An alternative would be to introduce a “blinded C” dialect that integrates blinding directly into the type system, similar to the approaches discussed for Rust in § 1. For example, a blinded keyword akin to const could be used to mark variables as a blinded counterpart of their basic type. While this would improve developer ergonomics and simplify function parameter annotations, it would require intrusive changes to the language and compiler. This could hinder maintainability and extensibility, especially for evolving platforms like CHERI.
@@
As we note in § 6.1, BLACKOUT, unlike previous data-oblivious ISAs (Yu et al., 2019a; ElAtali et al., 2024b), does not add any additional instructions to the underlying ISA but leverages existing CHERI functionality with only small adjustments to its permission model to facilitate necessary architectural changes to support blinded capabilities. Future work can extend the ISA by adding dedicated instructions to support blinded memory. A potential extension is an instruction to directly set the blindedness bit in a register without a load via a blinded capability. This would allow the blinded capability-enhanced compiler to optimize code where a certain variable can be kept completely in a register throughout its lifetime. Currently, BLACKOUT requires such variables to be allocated on the stack in order to load them via the corresponding blinded capability.
@@
BLACKOUT demonstrates a practical method for enhancing CHERI’s security model without invasive architectural changes by integrating data-oblivious computation capabilities into the existing CHERI protection model. This integration provides insights into how capability-based architectures can evolve to support advanced security properties, such as data-oblivious computation, while maintaining compatibility and performance. Consequently, BLACKOUT’s principles and mechanisms can inform ongoing CHERI standardization efforts (CheriAlliance, 2025), potentially guiding the evolution of CHERI-enabled architectures towards broader security guarantees. Above, we discussed adding custom instructions. But maintaining the CHERI ISA as we currently do in BLACKOUT provides backward-compatibility for non-BLACKOUT hardware and can ease integration into the standard.
Relationship to non-interference.
Since BLACKOUT builds on BliMe, it also ensures (as shown in Section VI of (ElAtali et al., 2024b)) the values of blinded data have no effect on the rest of the system by 1) preventing blinded values from directly flowing to insecure instructions, and 2) by ensuring no secret-dependent branches are allowed when operating on blinded data, with an exception for controlled declassification as described above and in § 5.2. Given its scope, we conjecture that this guarantee satisfies the non-interference property even though it was not explicitly discussed in (ElAtali et al., 2024b). Providing a formal model for BLACKOUT itself and explicitly establishing non-interference are left for future work. Nevertheless, we empirically investigate non-interference for BLACKOUT in § 7.3.